# Polynomial Composition Activations: Unleashing the Dynamics of Large Language Models

### v1

@@ Appendix D Additional Results on Dense Model
More detailed results from our ablation studies are shown in Figures 8, 9, and 10. These figures illustrate the training loss, validation loss, and validation perplexity (PPL) for the 1B dense model under different configurations.
@@
Figure 8: Traning loss, validation loss, and validation perplexity (PPL) for the 1B dense model with different orders of PolyReLU activation functions.
@@
Figure 9: Traning loss, validation loss, and validation perplexity (PPL) for the 1B dense model with different polynomial compositions.
@@
Figure 10: Traning loss, validation loss, and validation perplexity (PPL) for the 1B dense model with different variants of ReLU activation functions.
Appendix E Additional Results on MoE model
@@
More results for the MoE model are provided in Figure 11, , showcasing validation losses and downstream evaluations after 200 billion training tokens. The comparison highlights models with different activation functions, such as SwiGLU and PolyNorm. As shown, models with PolyNorm exhibit lower training and validation losses, along with superior downstream performance.

### v2

@@ Appendix I Additional Results on Dense Model
More detailed results from our ablation studies are shown in Figures 10, 11, and 12. These figures illustrate the training loss, validation loss, and validation perplexity (PPL) for the 1B dense model under different configurations.
The results of the 1B dense models trained on 400 billion tokens are presented in Figure 13. As shown in the figure, models employing PolyReLU and PolyNorm consistently achieve significantly better performance compared to SwiGLU.
@@
Figure 10: training loss, validation loss, and validation perplexity (PPL) for the 1B dense model with different orders of PolyReLU activation functions.
@@
Figure 11: training loss, validation loss, and validation perplexity (PPL) for the 1B dense model with different polynomial compositions.
@@
Figure 12: Training loss, validation loss, and validation perplexity (PPL) for the 1B dense model with different variants of ReLU activation functions.
Figure 13: Training loss, validation loss, and validation perplexity (PPL) for the 1B dense model with 400 billion training tokens.
Appendix J Additional Results on MoE model
@@
More results for the MoE model are provided in Figure 14, showcasing validation losses and downstream evaluations after 200 billion training tokens. The comparison highlights models with different activation functions, such as SwiGLU and PolyNorm. As shown, models with PolyNorm exhibit lower training and validation losses, along with superior downstream performance.